Circular No. 7617

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1118+480 J. Casares, C. Zurita, T. Shahbaz, and P. Rodriguez-Gil, Instituto

de Astrofíisica de Canarias (IAC); P. Charles and R. Hynes, University of Southampton; R. M. Wagner, LBT Observatory; E. Ryan, University of Arizona; C. Foltz, MMT Observatory; and S. Starrfield, Arizona State University, report on the preliminary analysis of long-term R-band CCD photometry of the blackhole x-ray transient XTE J1118+480 (cf. IAUC 7389): “Observations obtained on 41 nights from Dec. 2000 to Apr. 2001 with the IAC-80 and 1-m Optical Ground Station telescopes at Tenerife, the 2.56-m Nordic Optical Telescope and 1-m Jacobus Kapteyn Telescope at La Palma, and the MDM 1.3-m telescope at Kitt Peak reveal that XTE J1118+480 is still fading from its early 2000 outburst (IAUC 7390) at a rate of ∼ 0.003 mag/day, and it is currently at R = 18.6 (2001 Apr. 25). After subtraction of a nightly mean magnitude, the Fourier power spectrum of these data is dominated by the orbital period, P_{orb}, with \frac{1}{2}P_{orb} = 0.084968(1) day. Additionally, large night-to-night variations (amplitude 0.08 mag) are clearly detected in the depth of the minima and the overall lightcurve shape, indicative of a short-timescale distorting wave. After subtraction of the orbital (ellipsoidal) modulation, a phase-dispersion-minimization analysis (cf. Stellingwerf 1978, Ap.J. 224, 953) reveals prominent peaks at a period that we intrepret as the superhump period (caused by a precessing eccentric disk), P_{sh} = 0.17049(1) day, as well as at \frac{1}{2}P_{sh} and \frac{1}{4}P_{sh}. With this P_{sh} (and consistent with that seen in outburst; cf. IAUC 7418), the disk precession period must therefore be ∼ 52 days, and the power spectrum of our original data indeed shows a low-frequency peak at ∼ 45 days. Extrapolating the superhump period-excess-versus-mass-ratio relation (Osaki 1985, A.Ap. 144, 376) for cataclysmic variables, our value of P_{sh} = 1.0033P_{orb} implies the most extreme mass ratio in this class with M_{2}/M_{1} ∼ 0.008; however, we note that for this extreme ratio, the 2:1 resonance rather than the classical 3:1 resonance might be the dominant tidal instability (Whitehurst and King 1991, MNRAS 249, 25). Our spectroscopic database (see Wag-

ner et al., http://xxx.lanl.gov/abs/astro-ph/?0104032) shows nightly

bulk motions in Hα of a few hundred km/s that may represent an elliptical precessing disk. Further spectroscopy and photometry is urged in order to monitor this behavior as the disk decays.”

2001 May 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green